# Adapting Large Language Models for Education: Foundational Capabilities, Potentials, and Challenges

### v2

@@ 2.3.2 Code Refinement
In most cases, LLMs could not generate the correct code at once. We could enable LLMs to generate a code sketch (either actual code or pseudocode) and utilize various methods to guide the model to modify and refine the code. By leveraging the inherent code correction ability of LLMs, the overall precision and quality of the code could be significantly enhanced. Liventsev et al. (2023) constructed a pipeline: Synthesize, Execute, Instruct, Debug, and Rank (SEIDR). It first generates multiple different codes and undergoes the process of code filtering and debugging, ultimately selecting the best code among them. According to Magister et al. (2022), teaching an LLM to debug its program draft via few-shot demonstrations could improve the performance on code generation tasks. Another method for LLM debug is generating unit tests by LLM itself and checking its code (Chen et al., 2022a). By mimicking the human coding process, LLM’s programming ability is greatly enhanced. However, these methods lead to an increased number of calls to LLMs, resulting in a significant increase in inference time.
@@
In coding education, the guidance given by current LLMs still falls short of the level offered by human instructors. Firstly, there is significant room for improvement in the coding abilities of LLMs. When it comes to generating complex algorithms, LLMs’ performance rapidly declines compared to the functionality achieved by humans (Chen et al., 2021b). Additionally, due to the lack of real-world data, LLMs struggle to learn the intermediate thinking process of code writing, making it difficult for them to provide relevant explanations and instructions to beginners. As a result, the use of LLMs in programming education still needs to be improved, especially in terms of interpretability.
2.4 Reasoning
@@
The reasoning capability of LLMs offers significant potential for educational use. LLMs can act as advanced tools, enhancing students’ cognitive processes, providing personalized mentorship, and offering tailored learning support. Consequently, these models can bolster students’ problem-solving skills and promote holistic cognitive development in an educational setting. This section reviews LLMs’ general reasoning ability development strategies.
2.4.1 Supervised Fine-tuning for Reasoning
Prior researches on the LLMs’ reasoning ability are primarily based on the fully supervised fine-tuning method. For instance, Rajani et al. (2019) effectively fine-tuned a pre-trained GPT model to generate rationales for predictions based on the CoS-E dataset (Talmor et al., 2018), enhancing performance in commonsense question-answering tasks. However, the success of fine-tuning mainly depends on the availability of a dataset with explicit reasoning, which could be challenging to obtain. Furthermore, the inferences drawn from fine-tuning are mainly limited to specific domain datasets, relying more on the content within the fine-tuning data than on the LLM’s inherent reasoning abilities.
2.4.2 Prompt Engineering for Reasoning
Efforts have been made in recent research to tackle the constraints inherent in the fine-tuning process of LLMs. These fine-tuning methods tend to overfit specific dataset distributions, reducing their effectiveness on more diverse datasets. To mitigate this issue, numerous studies have developed strategies that utilize the robust reasoning abilities of LLMs, taking advantage of their extensive pre-trained parameters. One such strategy involves guiding LLMs to generate inference and reasoning processes through demonstrations or prompts. For example, Wei et al. (2022) introduced the ”Chain of Thought” (CoT) method, which employed natural language reasoning steps as prompts to enhance the reasoning performance of LLMs. Similarly, Wang et al. (2022) proposed a self-consistency strategy involving sampling a variety of reasoning paths and selecting the most consistent answer, which has been shown to improve model performance. Addressing the constraint of fixed manual annotation demonstrations, Diao et al. (2023) introduced an active selection approach. This method dynamically identified the most effective demonstrations relevant to task requirements from a pool of queries. Furthermore, Zhou et al. (2022) proposed a prompting strategy that decomposed complex problems into simpler sub-problems. This approach encouraged sequential problem-solving and could potentially enhance the performance of LLMs further.
2.4.3 Hybrid Methodologies for Reasoning
Despite the success of prompt engineering in improving the performance of LLMs, it does not boost their core reasoning skills by not changing the model parameters. To address this limitation, new approaches combining fine-tuning with prompt engineering have been developed, aiming to enhance the model’s reasoning abilities and responsiveness to prompts. One approach employs LLMs in a pedagogical capacity to instruct smaller student models. Studies by Ho et al. (2022) and Magister et al. (2022) demonstrated significant performance gains across various datasets through fine-tuning student models based on the reasoning outputs of larger LLMs, even for smaller-scale models. Moreover, Zelikman et al. (2022) reported substantial performance improvements across multiple datasets by generating step-by-step rationales and fine-tuning models based on correct answers, thus facilitating model learning from its reasoning.
@@ 2.4.3 Hybrid Methodologies for Reasoning
In conclusion, the discussed methods have significantly advanced the reasoning capability of LLMs, optimally utilizing their unique features for diverse reasoning tasks. This enhancement can greatly benefit educational applications. However, it’s crucial to recognize the limitations. As underscored by Valmeekam et al. (2022) and Ruis et al. (2022), LLMs struggle with complex reasoning tasks and those requiring implicit expressions. For example, LLMs can struggle with complex reasoning scenarios, leading to a notable decrease in performance. This is particularly relevant in educational contexts, where incorrect problem-solving modeled by LLMs could misguide students and lead to misunderstandings or flawed comprehension. Thus, despite LLMs’ immense potential in education, their limitations must be carefully considered to ensure they facilitate rather than obstruct learning.

### v3

@@ 3.3.2. Code Refinement
We investigate the development of LLMs in code refinement from two perspectives. One aspect is the advancement of LLMs in the task of fixing code bugs, known as Automated Program Repair (APR). Sobania et al. (2023) conducted experiments analyzing the performance of ChatGPT on the APR task. They found that it can achieve competitive results compared to previous deep learning-based methods, and incorporating additional information through dialogue can surpass previous approaches. Xia and Zhang (2023) introduced conversational APR, enabling LLMs to obtain bug feedback through dialogue, effectively enhancing the performance of various LLMs in APR. In addition to the APR task itself, the second aspect involves integrating code refinement into the process of code generation, leveraging bug feedback to enhance the effectiveness of code generation. Liventsev et al. (2023) constructed a pipeline: Synthesize, Execute, Instruct, Debug, and Rank (SEIDR). It first generates multiple different codes and undergoes the process of code filtering and debugging, ultimately selecting the best code among them. According to Magister et al. (2022), teaching an LLM to debug its program draft via few-shot demonstrations could improve the performance on code generation tasks. Another method for LLM debug is generating unit tests by LLM itself and checking its code (Chen et al., 2022b). By mimicking the human coding process, LLM’s programming ability is greatly enhanced. However, these methods lead to an increased number of calls to LLMs, resulting in a significant increase in inference time.
@@
In the context of coding education, the support and guidance provided by LLMs do not yet match the level of assistance offered by human instructors. One major reason for this gap is that LLMs still have substantial room for improvement in their coding capabilities. While LLMs can generate functional code for relatively simple tasks, when it comes to generating complex algorithms, LLMs’ performance rapidly declines compared to the functionality achieved by humans (Chen et al., 2021b). Additionally, due to the lack of real-world data, LLMs struggle to learn the intermediate thinking process of code writing, making it difficult for them to provide relevant explanations and instructions to beginners. As a result, the use of LLMs in programming education still needs to be improved, especially in terms of interpretability.
3.3.3. Summary
Code data is more abundant in the training corpora of LLMs compared to mathematical data. This is primarily due to the inherent nature of coding, which heavily relies on computers and the internet. As LLMs are trained using data scraped from the web, they are exposed to a significant amount of code-related information during the training process. Many improvements to LLMs’ code generation abilities are inspired by the human programming process. For instance, programmers often refer to documentation and resources to gain a better understanding of the problem and potential solutions. For complex coding tasks, the solving process typically involves a cycle of design, writing, and debugging. These thought processes can be utilized to enhance the programming effectiveness of LLMs. From an educational perspective, it is crucial for LLMs to not only generate correct code but also possess the ability to analyze and provide feedback on code written by students. This involves identifying issues, suggesting improvements, and offering explanations to help students learn and grow as programmers.
3.4. Reasoning
@@
The reasoning capability of LLMs offers significant potential for educational use, serving as advanced tools that enhance students’ cognitive processes, provide personalized mentorship, and offer tailored learning support. This section reviews LLMs’ general reasoning ability development strategies.
3.4.1. Supervised Fine-tuning for Reasoning
Previous studies have primarily focused on fully supervised fine-tuning LLMs to enhance their reasoning capabilities. This approach aligns model outputs closely with labeled datasets, allowing the models to produce highly accurate predictions within specific domains. One such study by (Rajani et al., 2019) demonstrated the efficacy of fine-tuning a pre-trained GPT model, which generated rationales for predictions on the CoS-E dataset (Talmor et al., 2018). The results revealed that models trained with explanations exhibited improved performance in commonsense question-answering tasks. However, the effectiveness of fine-tuning methods heavily relies on the availability of a specific dataset that includes explicit reasoning steps. Acquiring such a dataset can prove to be challenging. Moreover, the scope of inference from fine-tuned models is restricted to the dataset’s domain, hinging largely on the data’s inferential quality. This constraint highlights the benefits and limitations of fully supervised fine-tuning, as it narrows the model’s reasoning abilities to the dataset’s specific domain. Consequently, it underscores the need to explore methods that harness LLMs’ intrinsic reasoning capabilities, potentially providing broader relevance and deeper insights beyond the limitations of domain-specific datasets.
3.4.2. Prompt Engineering for Reasoning
Efforts have been made in recent research to tackle the constraints inherent in the fine-tuning process of LLMs. These fine-tuning methods tend to overfit specific dataset distributions, reducing their effectiveness on more diverse datasets. In response to this issue, a variety of strategies have been proposed. These strategies are designed to draw upon the robust reasoning abilities inherent in LLMs by leveraging their extensively pre-trained parameters. One approach involves guiding LLMs to generate inference and reasoning through demonstrations or prompts. For example, Wei et al. (2022) introduced the “Chain of Thought” (CoT) method, which utilized natural language reasoning steps as prompts for the model. By integrating CoT within a few-shot prompting framework, the model leveraged its extensive parameters to produce analogous chains of reasoning. Consequently, this approach empowered the model to adeptly navigate complex reasoning tasks across diverse domains, obviating additional training or fine-tuning. This innovation underscored the model’s inherent capability to generate deductive pathways, significantly enhancing its applicability and versatility in problem-solving scenarios without extensive domain-specific adaptations. Similarly, Wang et al. (2022) introduced a self-consistency strategy that enhances model performance by sampling various reasoning paths and selecting the most consistent answer. This approach diversified the exploration of reasoning strategies. It ensured the conclusions’ reliability, showcasing an innovative way to leverage the model’s capabilities for improved decision-making and problem-solving across different contexts. Confronting the limitations of relying on static, manually annotated demonstrations, which can restrict the adaptability of LLMs to the varying complexities of real-world tasks, Diao et al. (2023) introduced an active selection approach. This technique dynamically pinpointed the most pertinent demonstrations aligned with the specific demands of a task from a broad set of queries. In this way, the approach enhanced the flexibility and effectiveness of LLMs in adapting to diverse and evolving problem contexts. Concurrently, Zhou et al. (2022b) devised a prompting methodology that broke down intricate problems into their simpler constituent sub-problems. This tactic not only promoted a step-by-step problem-solving process but also hold promise for augmenting the efficacy of LLMs in handling complex tasks.
Building upon the CoT methodology, subsequent developments have introduced more sophisticated frameworks for enhancing the reasoning capabilities of LLMs. The Tree of Thoughts (ToT) (Yao et al., 2024) framework extended CoT by enabling LLMs to explore multiple reasoning paths through a hierarchical structure, thereby improving decision-making for tasks requiring strategic planning. Following the ToT, the Boosting of Thoughts (BoT) (Chen et al., 2024) framework introduced a novel approach by iteratively exploring and self-evaluating multiple trees of thoughts. This process accumulated an ensemble of trial-and-error reasoning experiences, offering a new form of prompting designed to tackle complex problems. Starting with simple prompts, BOT iteratively refined reasoning steps through error analysis, significantly improving the generation of reasoning paths and achieving higher problem-solving rates on complex tasks than existing advanced prompting strategies. To structure thoughts through prompts without depending on fine-tuning, the Graph of Thoughts (GoT) framework (Besta et al., 2023) introduced a new angle by arranging thoughts generated by LLMs into a graph structure. This setup fostered a dynamic interaction among different thought units, facilitating synthesizing synergistic results, simplifying intricate thought networks, and refining ideas via feedback mechanisms. GoT’s graph-based approach presented a versatile tool for problem-solving, allowing for a more nuanced and interconnected reasoning process that mirrors the complexity of human thought.
The emergence of CoT and related prompting strategies marked a significant evolution in utilizing LLMs for advanced reasoning, transitioning away from dependence on fine-tuning. These methods exploited the LLMs’ inherent capabilities to enhance their flexibility and effectiveness across various tasks without reliance on domain-specific tuning.
3.4.3. Hybrid Methodologies for Reasoning
Despite the success of prompt engineering in leveraging the intrinsic characteristics and capabilities of LLMs to enhance their performance, this approach falls short of fundamentally augmenting the model’s core reasoning abilities, as it does not alter the model’s underlying parameters. This inherent limitation points to a need for strategies that not only exploit the pre-existing strengths of LLMs but also seek to expand their innate capabilities. Innovative approaches that integrate the specificity of fine-tuning with the flexibility of prompt engineering have been developed to bridge this gap. These hybrid methodologies aim to bolster the LLMs’ responsiveness to complex prompts and substantially improve their intrinsic reasoning capacities, offering a more comprehensive enhancement of their problem-solving prowess. One practical approach is to utilize LLMs to “teach” language models with smaller model sizes. Ho et al. (2022); Magister et al. (2022) explored to fine-tune a student model on the chain of thought outputs generated by a larger teacher model and proved that enriching the fine-tuning data with such diverse reasoning results in a substantial performance boost across datasets even for very small models. Moreover, Zelikman et al. (2022) reported significant performance improvements across multiple datasets by generating step-by-step rationales and fine-tuning models based on correct answers, thus facilitating model learning from its reasoning. Similarly, Huang et al. (2022) proposed that by employing chain-of-thought prompting (Wei et al., 2022) and self-consistency (Wang et al., 2022) to generate rationale-augmented answers, and then used these answers for fine-tuning, LLMs autonomously refined their reasoning capabilities. This approach highlighted the significant ability of LLMs to advance their knowledge and problem-solving skills independently.
3.4.4. Summary
@@
In the educational domain, reasoning tasks possess unique characteristics that necessitate not only the accurate processing of information but also the capability to navigate and elucidate complex concepts in a manner that is accessible and educational for learners. As mentioned above, the discussed methods have significantly advanced the reasoning capability of LLMs, optimally utilizing their unique features for diverse reasoning tasks. This enhancement can greatly benefit educational applications. However, it’s crucial to recognize the limitations. As underscored by Valmeekam et al. (2022) and Ruis et al. (2022), LLMs struggle with complex reasoning tasks and those requiring implicit expressions. For example, LLMs can struggle with complex reasoning scenarios, leading to a notable decrease in performance. This is particularly relevant in educational contexts, where incorrect problem-solving modeled by LLMs could misguide students and lead to misunderstandings or flawed comprehension. Thus, despite LLMs’ immense educational potential, their limitations must be carefully considered to ensure they facilitate rather than obstruct learning.
3.5. Knowledge-based Question Answering
In the context of Knowledge-based Question Answering using LLM, the user presents a question to LLM, and LLM leverages knowledge-based methods and responds with the corresponding answer. Previous work by (Ren et al., 2023) showed that LLMs have an inaccurate perception of factual boundaries and often exhibit overconfidence. Many studies have explored and utilized external knowledge from open-world and domain-specific databases to enhance the knowledge base of these LLMs.
3.5.1. Open-domain QA
Open-domain question answering requires LLMs to accurately determine the reliability of information in the open world and craft their responses based on that understanding. The critical requirements for open-domain QA are real-time responsiveness and authenticity. LLMs exhibit disadvantages in both of these aspects. Due to the fixed parameters of the model, ensuring real-time information solely through the LLM itself is challenging, and LLMs commonly suffer from severe hallucination issues (Xu et al., 2024), posing a challenge to their authenticity as well. Jiang et al. (2021) evaluated the accuracy of LLM responses to a particular question from the perspective of calibration. Through experiments, the researchers discovered that models such as T5, BERT, and GPT-2 are not well-calibrated in QA tasks. While suggesting that incorporating calibration-related methods into the fine-tuning process can effectively enhance performance in QA tasks, it is evident that solely pre-trained language models still face significant challenges in open-domain tasks. To overcome this challenge, many works tried to add additional information to help the LLMs answer correctly (Khandelwal et al., 2019; Guu et al., 2020; Borgeaud et al., 2022). A common information source used is from the web. Lazaridou et al. (2022) employed information gathered from web searches as prompt input for LLMs, conditioning it to generate answers to questions. This approach effectively enables LLMs to use open-world information to answer questions.  Vu et al. (2023) introduced FreshPrompt, which incorporated web pages collected from the internet into prompts given to large-scale models. This allowed them to leverage the latest information when answering questions.  Kasai et al. (2024) developed a QA platform REALTIME QA that updated itself weekly. Through evaluation on this platform, they found that GPT-3 could update its generation results based on newly retrieved documents. However, when retrieved documents fail to provide sufficient information to find the answer, GPT-3 may provide outdated answers.
3.5.2. Domain-specific QA
Although LLMs are trained on vast corpora, they may still exhibit gaps in understanding specific domains. The primary challenge faced by LLMs in this task is the lack of domain knowledge. For specific domain questions, providing good answers often requires a considerable amount of expertise or skills in that field, while professional data is relatively scarce in the corpus of large-scale models. One straightforward approach is to fine-tune LLMs on specialized datasets. Typically, there are dedicated knowledge repositories for professional content that consolidate domain-specific knowledge, such as MedlinePlus [^1] https://medlineplus.gov, GeeksforGeeks [^2] https://www.geeksforgeeks.org, etc. Choi et al. (2023) utilized an external knowledge base to generate a set of question-answer pairs and then employed fine-tuning to transfer financial knowledge to LLMs, significantly improving financial question-answering tasks. Another common approach is to leverage the in-context learning capability of LLMs by incorporating retrieved knowledge from the knowledge base into prompts. Peng et al. (2023b) demonstrate this approach in their work on pest identification. They first use text embeddings, which are dense vector representations of text, to retrieve relevant information from a knowledge base. Text embeddings allow for efficient and accurate retrieval of similar or related content based on the semantic similarity between the query and the stored information. Once the relevant knowledge is retrieved, it is incorporated into the prompts provided to the LLMs. The LLMs then utilize their automatic feature extraction capabilities to process and understand the retrieved information in the context of the pest identification task. zhang et al. (2023) utilized K-nearest neighbors (KNN) (Guo et al., 2003) to search for the most similar K records from an accounting database, serving as k-shot examples, and greatly improved accounting efficiency. There are also works that train and improve the retriever encoder (Zhang et al., 2023b), as well as distill and refine the data in the database (Jeronymo et al., 2023). Such retrieval frameworks have lower costs and can be more flexible in applications across different domains. Liu et al. (2023d) proposed the RETA-LLM, a system that leveraged an information retrieval system based on Google Search to initially retrieve the top-k documents relevant to a user’s query, allowing LLMs to generate answers based on these retrieved documents. Furthermore, the system included plug-and-play modules that enable users to construct their own domain-specific LLMs. These modules covered various functionalities, including request rewriting, document retrieval, passage extraction, answer generation, and fact-checking.
3.5.3. Summary
While LLMs have mastered a broad range of open-world knowledge through extensive corpus training, their fixed parameters make it challenging to handle real-time, high-demand open-domain questions. Severe hallucination issues further compromise their accuracy in both open-domain and domain-specific queries. In the field of education, where authenticity is paramount, students may pose questions about textbook knowledge points. If the accuracy of these responses cannot be guaranteed, it could potentially mislead students. Therefore, a feasible solution to the inherent hallucination issues in foundational LLMs is to integrate external information, such as authoritative documents, allowing LLMs to base their responses on such external sources to mitigate hallucination problems.
Figure 3. A summary framework diagram for the approaches of LLMs in the development of education-related abilities. It categorizes previous enhancement strategies into three parts: Input Data Refinement, Model Self-Improvement, and External Tool Usage.
3.6. Discussion
Despite the varied contexts and specific challenges associated with each capability, certain strategies and insights resonate universally among researchers working to harness the capabilities of LLMs for educational purposes. Here we discuss the trends or commonalities we discovered across the development of capabilities.
3.6.1. High-quality data could help LLMs develop capabilities effectively.
Since the advent of the deep learning era, high-quality training data has significantly improved model performance. This is especially true for LLMs, and the approach extends beyond mere training. Through in-context learning, where desired inputs and outputs are formed into demonstrations and provided as prompts to LLMs, if the demonstrations are well-chosen, they can also greatly enhance the model’s capabilities.
There is a notable trend towards leveraging smaller models fine-tuned on high-quality data to surpass the performance of larger language models. This approach emphasizes the importance of focused, domain-specific training rather than simply relying on the vast number of parameters in large-scale models. The term “high-quality data” here refers to data that provides detailed supervisory signals for specific domain problems. For instance, even powerful LLMs like GPT-4 encounter high error rates with large number multiplication in basic arithmetic problems. However, with a dataset that includes detailed steps for multiplication and addition, even a small transformer model can effectively solve these problems. Similarly, for reasoning tasks, it’s possible to improve LLMs’ performance on these issues without altering model parameters, simply by incorporating specific reasoning steps into the demonstrations in the prompt. This highlights that the application of “high-quality data” in LLMs extends beyond just fine-tuning.
There is a concerted effort among researchers to explore how smaller models can achieve equivalent or superior outcomes in certain scenarios where LLMs have already exceeded human benchmarks. This pursuit reflects a broader shift towards optimizing computational efficiency and model scalability, ensuring that the advancements in LLM technology remain accessible and sustainable. The primary scenarios and objectives for training small models can be divided into two categories:
  • Training specialized models with high-quality data. For applications within specific, narrow fields, such as programming, developing a small, specialized model through data collection can facilitate deployment and reduce the computational resources needed for inference, among other benefits.
  • Training small models through knowledge distillation. In cases where the required knowledge and skills for an application are more general, it may be challenging to construct a specific dataset for training a small model. By learning to match the outputs of the teacher model, the student model can effectively absorb the knowledge and skills of the larger model, without the need for a specific, curated dataset. This allows the smaller model to inherit the generalization capabilities of the large teacher model, while being more computationally efficient and easier to deploy in resource-constrained environments.
Efficiency is an unavoidable issue for LLMs during training and real-world deployment. Training-wise, low-loss, high-efficiency training schemes like LoRA (Low-Rank Adaptation) (Hu et al., 2021) are continually being introduced. These can significantly reduce the number of trainable parameters required during fine-tuning. However, small models are still needed. In practical applications, where the trained models need to be deployed and used for generating predictions or outputs, the size of the model still plays a critical role.
3.6.2. LLMs could achieve self-improvement.
The inference abilities and text comprehension skills of LLMs enable them to conveniently obtain feedback, which allows them to refine their outputs. This process is known as the self-improvement, which is a general approach in improving LLMs’ answers to all the capabilities. LLMs can achieve self-improvement through a methodical approach that involves iterative refinement and multiple sampling. For instance, to provide better responses to queries, an LLM may first generate an initial output. Then, it evaluates the output’s effectiveness or accuracy. Leveraging multiple sampling, the LLM explores different solution pathways or creative responses, which expands its potential answers pool. Through iterative refinement, it compares, contrasts, and consolidates these possibilities, learning which strategies yield the best results. This could involve internal processes such as adjusting parameters based on feedback loops, where it might integrate data from new examples or corrections provided by human users. Over time, this enhances the LLM’s ability to provide more precise, informative, and contextually relevant answers, thus gradually improving its problem-solving and content creation skills. In addition to individual LLMs enhancing their capabilities through feedback, the method of coordinating multiple LLMs to improve overall output is also being explored across various fields. For example, in solving programming problems, the process can be segmented into different stages, with each stage managed by a distinct LLM. These LLMs communicate and collaborate to complete the task collectively. Alternatively, one LLM may act as a generator to produce answers, while another serves as an evaluator to provide feedback. Through continuous dialogue between the two, the response can be consistently assessed and refined, thereby improving the quality of the answer. This cooperative approach leverages the strengths of different models to achieve a more effective and sophisticated problem-solving mechanism. Although the method of self-improvement is effective, it often results in a longer time to produce responses. For an LLM-based education system, the importance of providing accurate answers to students’ questions (to avoid misleading them) outweighs the need for speed. Therefore, employing a multi-agent approach to enhance answer quality through LLMs’ collaboration, or using sampling and iterative optimization for self-improvement, is an appropriate strategy for developing an education system. This ensures that the system prioritizes the correctness and reliability of information, which is crucial in educational settings where students’ learning effect is depend on the accuracy of the content provided.
3.6.3. Calling external tools is an universal method.
The integration of external tools into the LLM framework is a widely adopted method. This strategy not only enhances the models’ ability to access and incorporate real-time information and authoritative sources but also mitigates some of the inherent limitations of LLMs, such as their tendency towards factual inaccuracies or hallucinations. We can divide the use of external tools by LLMs into two perspectives:
  • LLMs inherently have certain limitations that cannot be resolved through training alone, and external tools can be used to address these deficiencies. In this scenario, tools serve the LLMs. For example, LLMs’ high error rates in large number multiplication can easily be mitigated by employing an external calculator. Similarly, LLMs’ inability to access real-time information can be compensated for by retrieving the latest web pages via web API calls.
  • Utilizing the reasoning and decision-making capabilities of LLMs, the invocation of external tools can influence the real world. In this approach, the primary task of LLMs is to make informed decisions about when and which tools to utilize in order to accomplish specific tasks. The key responsibility of LLMs in this approach is to act as intelligent agents that can analyze a given situation, understand the requirements and constraints, and determine the most appropriate course of action.
The LLM that solves problems by invoking external tools is a type of LLM agent, where the external tool is not necessarily an API but can also be an expert model. Fine-tuning LLMs on specific datasets can yield excellent results on the corresponding task, but it’s impractical to fine-tune LLMs across datasets for all capabilities. A viable solution is to use fine-tuned small language models as expert models, which serve as external tools for the central LLM. Compared to APIs, the advantage of trained expert language models is their ability to understand more granular and flexible demands from the LLM, providing targeted feedback.